Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION)

9 sin 2θ − 2 sin θ = 0

Answer by lwsshak3(11628)

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION)
For: [0,90˚]
9 sin 2θ − 2 sin θ = 0
9(2sin θcos θ)-2sin θ=0
sin θ(18cos θ-2)=0
sin θ=0
θ=0
or
18cos θ-2=0
cosθ=2/18=1/9
θ=83.621˚